# Eigenspectrum Analysis of Neural Networks without Aspect Ratio Bias

### v1

@@ 3.2 Rationale of FARMS: Why Typical HT-SR Methods Are Insufficient
Here, we explain the rationale behind our method FARMS. In Section 3.1, we mentioned that layers exhibiting more HTed ESDs tend to be more well-trained. This observation is the key to using HT-SR methods to measure model quality. However, beyond the training quality of weights, the aspect ratio of a weight matrix also influences the heavytailness of its ESD. Specifically, in RMT, the Marchenko-Pastur (MP) distribution describes the limiting behavior of singular values in large rectangular random matrices. According to the MP distribution, the ESD of the correlation matrix $\mathbf{W}^{\top}\mathbf{W}$ of an i.i.d. Gaussian random matrix $\mathbf{W}_{m\times n}$ becomes more concentrated as the aspect ratio $m/n$ deviates from 1. An example is presented in Figure 2. Consequently, weight matrices with different aspect ratios naturally exhibit varying ESD shapes, which can interfere with the quality estimation of model layers. We refer to this issue as the aspect ratio bias in HT-SR methods.
Figure 2: ESD shapes can be biased by the aspect ratio $m/n$. The left ESD is more HTed, while the right ESD is more concentrated. However, this change in the shape of the ESD is entirely due to the different aspect ratios of these two matrices and is not caused by differences in training quality. See detailed settings and supplementary results in Appendix B Figure 8.
@@ 3.3 Analyzing ESDs Using FARMS
Subsequently, for 2D Linear layers we compute the eigenvalues of the correlation matrices $\mathbf{X}_{ij}=\mathbf{W}_{ij}^{\top}\mathbf{W}_{ij}$ for each submatrix in the $i$-th layer. We then average the ESDs of $\mathbf{W}_{i1},\mathbf{W}_{i2},\dots,\mathbf{W}_{il}$ (by concatenating their corresponding eigenvalue series) and measure the HT metrics of the averaged ESD instead. We concatenate the eigenvalue series because that’s equivalent to averaging the empirical densities. The average of the ESDs leads to a less variant estimate of HT characteristics.
For CNN layers, since they have 4D tensor weight matrices with four dimensions $[C_{1},C_{2},k_{H},k_{W}]$ (which represent input channels, output channels, height, and width, respectively), we employ a slightly different method for subsampling and measuring. We first flatten the two dimensions representing the convolution kernel size, $[k_{H},k_{W}]$, in the 4D tensor into a single dimension. This results in $l^{\prime}=k_{H}\times k_{W}$ two-dimensional matrices of shape $[C_{1},C_{2}]$. We assume here that $C_{1}\geq C_{2}$ for convenience. For a 3D tensor of shape $l^{\prime}\times C_{1}\times C_{2}$, we perform subsampling of size $m^{\prime}\times n^{\prime}$ on each $C_{1}\times C_{2}$ matrix corresponding to each $l^{\prime}$. This results in $[\frac{C_{1}}{m^{\prime}}]\times[\frac{C_{2}}{n^{\prime}}]\times l^{\prime}$ submatrices, each of size $m^{\prime}\times n^{\prime}$. Next, we average the ESDs by concatenating the rooted singular values for each submatrix along the $l^{\prime}$ dimension, and then measure the HT metrics of the ESDs. In this way, we obtain $[\frac{C_{1}}{m^{\prime}}]\times[\frac{C_{2}}{n^{\prime}}]$ HT metrics values. Finally, we take the average of these metrics values to obtain the HT metrics corresponding to the 2D CNN. We also considered concatenating all the rooted singular values into a single list and calculating the corresponding ESD to measure the degree of heavy-tailedness. However, our experimental results in Table  7 in Appendix  C.2 show that averaging the results separately leads to better model performance. The detailed procedures of FARMS are shown in Figure 1 and Figure  3.
Figure 3: Main Steps in FARMS.
@@ 4 Empirical Results
4.1 Experimental Setup
Datasets. For image classification, we consider the CIFAR100 dataset (Krizhevsky et al., 2009). CIFAR100 consists of 50K pictures for the training set and 10K pictures for the testing set with 100 categories. For evaluating LLM pruning methods, we calculate model perplexity on the held-out WikiText (Merity et al., 2016) validation set and use seven tasks, including BoolQ (Clark et al., 2019), RTE (Wang, 2018), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC Easy and Challenge (Clark et al., 2018) and OpenbookQA (Mihaylov et al., 2018) for downstream zero-shot evaluation (Gao et al., 2021). For SciML, we fine-tune the models on simulated solutions of time-dependent PDE dataset 2D Compressible Navier-Stokes (CFD)  [^3] CFD means compressible fluid dynamics or, equivalently, the compressible Navier-Stokes equations. from PDEBench (Takamoto et al., 2022). All datasets considered in this paper are standard and widely studied.
Models. We consider different types of NNs in various research fields: VGG (Simonyan and Zisserman, 2014), ResNet (He et al., 2016), OPT (Zhang et al., 2022), LLaMA (Touvron et al., 2023a, b; Grattafiori et al., 2024) and DPOT (Hao et al., 2024). For VGG series, we consider VGG16 and VGG19. For ResNet series, we consider ResNet18 and ResNet34. For OPT series, we consider four different model size: OPT-125M/350M/1.3B/6.7B. For LLaMA series, we consider six different moodels: LLaMA-7B/13B, LLaMA-V2-7B/13B and LLaMA-V3-3B/8B. For DPOT series, we consider DPOT-Tiny and DPOT-Small.
@@ 4.1 Experimental Setup
0.7 AlphaPruning 231.76 24.30$\pm$ 0.25 18.66$\pm$ 0.49 2029.20 14.47$\pm$ 0.08 13.29$\pm$ 0.17
Ours 173.49 22.61$\pm$ 0.18 18.53$\pm$ 0.40 413.76 14.20$\pm$ 0.09 13.06$\pm$ 0.14
0.75 AlphaPruning 2046.22 104.53$\pm$4.49 36.52$\pm$1.13 2710.49 32.18$\pm$0.31 22.26$\pm$0.59
Ours 1704.56 71.67$\pm$1.71 35.47$\pm$1.01 2634.82 29.56$\pm$0.33 20.80$\pm$0.43
0.8 AlphaPruning 28865.67 772.20$\pm$78.70 96.02$\pm$1.59 5399.87 160.59$\pm$4.05 47.57$\pm$2.64
Ours 12799.58 504.58$\pm$23.05 79.42$\pm$3.86 5026.86 127.49$\pm$2.12 41.44$\pm$1.58
0.85 AlphaPruning 71710.96 4609.70$\pm$978.39 272.84$\pm$30.84 38140.95 3144.01$\pm$597.79 122.38$\pm$8.88
Ours 66808.51 3595.54$\pm$810.54 234.46$\pm$16.42 37453.06 2847.85$\pm$368.10 101.06$\pm$4.48
@@ 4.2 Improving LLM Pruning with FARMS
To explore the effectiveness of FARMS, we conduct experiments in the field of LLM pruning. As we have mentioned, we can replace the HT measurement procedures in AlphaPruning with FARMS. Therefore, we can compare it with FARMS and without. We make the comparison with different sparsity ratios in the range {0.7, 0.75, 0.8, 0.85} and different LLM pruning methods, including Magnitude-based (Han et al., 2015), SparseGPT (Frantar and Alistarh, 2023) and Wanda (Sun et al., 2023). We use different LLM pruning methods because AlphaPruning assigns layer-wise ratios but not precise pruning locations. We followed the experimental setup from Lu et al. (2024) and our hyperparameter ranges are reported in Table 14 of Appendix F.
@@
The results in Table 1 illustrate that using FARMS helps AlphaPruning achieve better performance under different sparsity ratios and LLM pruning methods. Specifically, when using the Magnitude method, our approach reduces the perplexity of LLaMA-7B from 231.76 to 173.49 at a sparsity ratio of 0.7. Similarly, when employing more advanced pruning methods such as Wanda and SparseGPT, our approach further reduces the perplexity of LLaMA-7B from 96.02 to 79.42 and LLaMA-13B from 47.57 to 41.44 at a sparsity ratio of 0.8. One notable advantage of AlphaPruning is that it allows pruning LLMs to a sparsity ratio of 0.8 without significantly impacting perplexity. With FARMS, this performance can be further improved. We can also find that FARMS achieves lower STD in most settings. In Table 8 and Table 9 in Appendix C, we provide additional experiment results on OPT series models and more recent LLaMA models.
@@ 4.3 Improving Image Classification with FARMS
In this subsection, we compare FARMS to TempBalance on image classification tasks. As we mentioned, TempBalance is a layer-wise learning rate assignment method using HT metrics. In Figure 4, we report the evaluation results of training the ResNet and VGG series models under different optimizing settings. Again, we compare TempBalance with or without FARMS. “Ours” means replacing the weight analysis method previously employed in the TempBalance with FARMS.
An important configuration when using the TempBalance method is the exclusion of certain layers from the learning rate assignment process. Instead of receiving layer-wise learning rate adjustments, these layers are assigned a global learning rate that decays according to cosine annealing. These excluded layers have “tall-and-skinny matrices,” which are matrices with a large aspect ratio and a relatively small number of eigenvalues. For example, the final layer of the ResNet 18 model has dimensions of $512\times 100$, resulting in an aspect ratio of $512/100$. This kind of extreme size makes it difficult for previous methods to accurately estimate the heavytailness of the ESDs. This issue further exacerbates the inability of the previous TempBalance method to accurately allocate per-layer learning rates, resulting in certain layers remaining at either excessively high or extremely low learning rates. We illustrate this phenomenon in more detail in Figure 12 of Appendix C. Consequently, this imbalance leads to suboptimal model performance after training.
Therefore, according to the experimental setup described in Zhou et al. (2024), the first and last layers of ResNet and VGG models, along with certain layers whose correlation matrix of weight matrix contains a relatively small number of eigenvalues, will be excluded from the layer-wise learning rate scheduling. To ensure a comprehensive study, we compare FARMS and TempBalance with and without this layer selection (LS) heuristic. See Figure 4, when the LS method is not applied (and all layers will be included in the adjustable learning rate scheduling), the performance of the original TempBalance method (shown as “TB (no LS)”) deteriorates significantly. For example, compared to the LS-enabled setup (shown as “LS+TB”), the test accuracy of the VGG 19 model trained on CIFAR100 without LS using the TempBalance method drops from 74.19% to 73.22%, while the standard deviation increases from 0.159 to 0.277. In contrast, when training models using the TempBalance method with FARMS, performance remains stable or even improves when LS is disabled. This observation demonstrates that our method is more robust to weight matrices of extreme sizes, and the reason is that our method provides a more accurate assessment of the training progress independent of the matrix size.
@@
TB 79.03$\pm$0.169 79.81$\pm$0.145 74.87$\pm$0.214 73.89$\pm$0.199
Ours 79.35$\pm$0.126 80.07$\pm$0.097 75.16$\pm$0.212 74.03$\pm$0.163
4.4 Improving SciML Models with FARMS
To explore the potential applications of FARMS in multiple domains of machine learning research, we also performed SciML fine-tuning experiments using the 2DCFD dataset with DPOT-Tiny and DPOT-Small models and compared FARMS with TB_Sigmoid method (Liu et al., 2024). We followed the experimental setup from Liu et al. (2024), and our hyperparameter ranges are detailed in Table 14 of Appendix F. In Table 4, we show the results of comparing FARMS with the TB_Sigmoid (shown as “TB_Sig”) with layer selection and the baseline fine-tuning (shown as “FT”) with a uniform learning rate for each layer in different data subsampling ratios in the range {5%, 10%, 25%, 50%, 100%}. Our method consistently outperformed TB_Sigmoid across all subsampling ratios. For instance, with a full dataset (100% of the data), our approach reduced the model’s L2RE to 1.017e-2. Comparing to the TB_Sigmoid, the L2RE decreased by 5.66%.
@@ 4.6 Ablation Study
Table 6: Different Window Size and Sampling Steps for pruning the LLaMA-7B at 0.8 sparsity ratio. The pruning method is SparseGPT. The model is evaluated on WikiText dataset with perplexity ()($\downarrow$)( ↓ ).
@@ 5 Conclusion
Impact Statement
This paper presents research aimed at advancing the fields of Machine Learning and Deep Learning, particularly in training using information from eigenspectrum analysis. While our work has various potential societal implications, we do not find it necessary to highlight any specific ones here.
Acknowledgments
This work is supported by DOE under Award Number DE-SC0025584 and Dartmouth College.
@@ Appendix B Marchenko–Pastur Distribution
Figure 8: The Marchenko-Pastur (MP) Law for different values of $n$. The ESD (blue histogram) of the eigenvalues of the sample covariance matrix is compared with the theoretical Marchenko-Pastur probability density function (red curve) for various aspect ratios $\gamma=m/n$, where $m=4000$ is fixed, and $n$ varies from 4000 to 16000.
@@
[…]
@@ C.2 Comparison of Different CNN Processing Methods
Calculating the Overall ESD 79.36$\pm$0.101 75.27$\pm$0.201
Averaging ESDs Computed from Subsets 79.53$\pm$0.177 75.36$\pm$0.118
@@ C.3 Experiment Results for Additional Models
We prune the OPT models  (Zhang et al., 2022) and LLaMA-V2/V3 models  (Touvron et al., 2023b; Grattafiori et al., 2024) with AlphaPruning and compare the model performance with FARMS or not. These models and their corresponding variants are widely used as benchmark models in LLM research (Yang et al., 2025b, a). We report the additional experiment results on LLM pruning in Table  8 and Table  9. For the OPT series models, we use four sparsity ratios: {0.6, 0.7, 0.75, 0.8}. For the LLaMA series models, we select {0.7, 0.75, 0.8, 0.85} as the sparsity ratios. Experimental results show that our method can help AlphaPruning achieve better model performance on models such as OPT.

### v2

@@ 3.2 Rationale of FARMS: Why Typical HT-SR Methods Are Insufficient
Here, we explain the rationale behind our method FARMS. In Section 3.1, we mentioned that layers exhibiting more HTed ESDs tend to be more well-trained. This observation is the key to using HT-SR methods to measure model quality. However, beyond the training quality of weights, the aspect ratio of a weight matrix also influences the heavytailness of its ESD. Specifically, in RMT, the Marchenko-Pastur (MP) distribution describes the limiting behavior of singular value distributions in large rectangular random matrices. According to the MP distribution, the ESD of the correlation matrix $\mathbf{W}^{\top}\mathbf{W}$ of an i.i.d. Gaussian random matrix $\mathbf{W}_{m\times n}$ becomes more concentrated as the aspect ratio $m/n$ deviates from 1. An example is presented in Figure 2. Consequently, weight matrices with different aspect ratios naturally exhibit varying ESD shapes, leading to different degrees of heavytailness, which, if not analyzed carefully, can interfere with the quality assessment of model layers. We refer to this issue as the aspect ratio bias in HT-SR methods.
Figure 2: ESD shapes can be biased by the aspect ratio. Here we visualize the eigenvalues of the symmetric matrix $\frac{1}{n}\mathbf{X}_{m}\mathbf{X}_{m}^{\top}$. The left ESD is more HTed, while the right ESD is more concentrated. However, this change in the shape of the ESD is entirely due to the different aspect ratios of these two matrices and is not caused by differences in training quality. See detailed settings and supplementary results in the Figure 8 in Appendix B.
@@ 3.3 Analyzing ESDs Using FARMS
Subsequently, for 2D Linear layers we compute the eigenvalues of the correlation matrices $\mathbf{X}_{ij}=\mathbf{W}_{ij}^{\top}\mathbf{W}_{ij}$ for each submatrix in the $i$-th layer. We then average the ESDs of $\mathbf{W}_{i1},\mathbf{W}_{i2},\dots,\mathbf{W}_{il}$ (by merging their corresponding eigenvalue series) and measure the HT metrics of the averaged ESD instead. We merge the eigenvalue series because that’s equivalent to averaging the empirical densities. The average of the ESDs leads to a less variant estimate of HT characteristics.
For CNN layers, since they have 4D tensor weight matrices with four dimensions $C_{1},C_{2},k_{H},k_{W}]$ (which represent input channels, output channels, height, and width, respectively), we employ a slightly different method for subsampling and measuring. We first flatten the two dimensions representing the convolution kernel size, $k_{H},k_{W}]$, in the 4D tensor into a single dimension. This results in $l^{\prime}=k_{H}\times k_{W}$ two-dimensional matrices of shape $C_{1},C_{2}]$. We assume here that $C_{1}\geq C_{2}$ for convenience. Then, for the 3D tensor of shape $l^{\prime}\times C_{1}\times C_{2}$, we perform subsampling of size $m^{\prime}\times n^{\prime}$ on each $C_{1}\times C_{2}$ matrix corresponding to each $l^{\prime}$. This results in $\frac{C_{1}}{m^{\prime}}]\times[\frac{C_{2}}{n^{\prime}}]\times l^{\prime}$ submatrices, each of size $m^{\prime}\times n^{\prime}$. Next, we average the ESDs by concatenating the rooted singular values for each submatrix along the $l^{\prime}$ dimension, and then measure the HT metrics of the ESDs. In this way, we obtain $\frac{C_{1}}{m^{\prime}}]\times[\frac{C_{2}}{n^{\prime}}]$ HT metrics values. Finally, we take the average of these metrics values to obtain the HT metrics corresponding to the original 2D CNN. We also considered concatenating all the rooted singular values into a single list and calculating the corresponding ESD to measure the degree of heavy-tailedness. However, our experimental results in Table 7 in Appendix C.2 show that averaging the results separately leads to better model performance. The detailed procedures of FARMS are shown in Figure 3.
@@ 4 Empirical Results
[…]
4.1 Experimental Setup
Datasets. For image classification, we consider the CIFAR100 dataset (Krizhevsky, 2012). CIFAR100 consists of 50K pictures for the training set and 10K pictures for the testing set with 100 categories. For evaluating LLM pruning methods, we calculate model perplexity on the held-out WikiText (Merity et al., 2017) validation set and use seven tasks, including BoolQ (Clark et al., 2019), RTE (Wang et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC Easy and Challenge (Clark et al., 2018) and OpenbookQA (Mihaylov et al., 2018) for downstream zero-shot evaluation (Gao et al., 2021). For SciML, we fine-tune the models on simulated solutions of time-dependent PDE dataset 2D Compressible Navier-Stokes (CFD)  [^3] CFD means compressible fluid dynamics or, equivalently, the compressible Navier-Stokes equations. from PDEBench (Takamoto et al., 2022). All datasets considered in this paper are standard and widely studied.
Models. We consider different types of NNs in various research fields: VGG (Simonyan and Zisserman, 2015), ResNet (He et al., 2016), OPT (Zhang et al., 2022), LLaMA (Touvron et al., 2023a, b; Grattafiori et al., 2024) and DPOT (Hao et al., 2024). For VGG series, we consider VGG16 and VGG19. For ResNet series, we consider ResNet18 and ResNet34. For OPT series, we consider four different model size: OPT-125M/350M/1.3B/6.7B. For LLaMA series, we consider six different moodels: LLaMA-7B/13B, LLaMA-V2-7B/13B and LLaMA-V3-3B/8B. For DPOT series, we consider DPOT-Tiny and DPOT-Small.
@@ 4.2 Improving LLM Pruning with FARMS
To explore the effectiveness of FARMS, we conduct experiments in the task of LLM pruning. As we have mentioned, we can replace the HT measurement procedures in AlphaPruning with FARMS. Therefore, we can compare AlphaPruning with FARMS and without. We make the comparison with different sparsity ratios in the range {0.7, 0.75, 0.8, 0.85} and different LLM pruning methods, including Magnitude-based (Han et al., 2015), SparseGPT (Frantar and Alistarh, 2023) and Wanda (Sun et al., 2024). We use different LLM pruning methods because AlphaPruning assigns only layer-wise ratios, not specific pruning locations, and thus can be applied to various LLM pruning methods. We followed the experimental setup from Lu et al. (2024) and our hyperparameter ranges are reported in Table 14 of Appendix F.
@@
The results in Table 1 illustrate that using FARMS helps AlphaPruning achieve better performance under different sparsity ratios and LLM pruning methods. Specifically, when using the Magnitude method, our approach reduces the perplexity of LLaMA-7B from 231.76 to 173.49 at a sparsity ratio of 0.7. Similarly, when employing more advanced pruning methods such as Wanda and SparseGPT, our approach further reduces the perplexity of LLaMA-7B from 96.02 to 79.42 and LLaMA-13B from 47.57 to 41.44 at a sparsity ratio of 0.8. One notable advantage of AlphaPruning is that it allows pruning LLMs to a sparsity ratio of 0.8 without significantly impacting perplexity. With FARMS, this performance can be further improved. We can also find that FARMS achieves lower perplexity STD in most settings. In Table 8 and Table 9 in Appendix C, we provide additional experiment results on OPT series models and more recent LLaMA models.
@@ 4.3 Improving Image Classification with FARMS
In this subsection, we compare FARMS to TempBalance on image classification tasks. As we mentioned, TempBalance is a layer-wise learning rate assignment method using HT metrics. In Figure 4, we report the evaluation results of training the ResNet and VGG series models under different optimizing settings. Again, we compare TempBalance with or without FARMS. “Ours” means replacing the method to measure heavytailness previously employed in the TempBalance with FARMS.
An important configuration used in the TempBalance method is the exclusion of certain layers from the learning rate assignment process. Instead of receiving layer-wise learning rate adjustments as in TempBalance, these layers are assigned a global learning rate that decays following the cosine annealing schedule. These excluded layers have “tall-and-skinny matrices,” which are matrices with a large aspect ratio and a relatively small number of eigenvalues. For example, the final layer of the ResNet 18 model has dimensions of $512\times 100$, resulting in a large aspect ratio of $512/100=5.12$. This kind of extreme size makes it difficult for previous methods to accurately estimate the heavytailness of the ESDs. This issue further exacerbates the inability of the previous TempBalance method to accurately allocate per-layer learning rates, resulting in certain layers remaining at either excessively high or extremely low learning rates. We illustrate this phenomenon in more detail in Figure 12 of Appendix C. Consequently, this imbalance leads to suboptimal model performance after training.
Therefore, according to the experimental setup described in TempBalance, the first and last layers of ResNet and VGG models, along with certain layers whose correlation matrix of weight matrix contains a relatively small number of eigenvalues, will be excluded from the layer-wise learning rate scheduling. To ensure a comprehensive study, we compare FARMS and TempBalance with and without this “layer selection” (LS) heuristic. See Figure 4, when the LS heuristic is not applied (and all layers will be included in the adjustable learning rate scheduling), the performance of the original TempBalance method (shown as “TB (no LS)”) deteriorates significantly. For example, compared to the LS-enabled setup (shown as “LS+TB”), the test accuracy of the VGG 19 model trained on CIFAR100 without LS using the TempBalance method drops from 74.19% to 73.22%, while the standard deviation increases from 0.159 to 0.277. In contrast, when training models using the TempBalance method with FARMS, performance remains stable or even improves when LS is disabled. This observation demonstrates that our method is more robust to weight matrices of extreme sizes, and the reason is that our method provides a more accurate assessment of the HTed ESDs independent of the matrix size.
@@
[…]
4.4 Improving SciML Fine-tuning with FARMS
To explore the potential applications of FARMS in multiple domains of machine learning research, we also performed SciML fine-tuning experiments using the 2DCFD dataset with DPOT-Tiny and DPOT-Small models and compared FARMS with the TB_Sigmoid method (Liu et al., 2024). We followed the experimental setup from Liu et al. (2024), and our hyperparameter ranges are detailed in Table 14 of Appendix F. In Table 4, we show the results of comparing FARMS with the TB_Sigmoid (shown as “TB_Sig”) with layer selection and the baseline fine-tuning (shown as “FT”) with a uniform learning rate for each layer in different data subsampling ratios in the range {5%, 10%, 25%, 50%, 100%}. Our method consistently outperformed TB_Sigmoid across all subsampling ratios. For instance, with a full dataset (100% of the data), our approach reduced the model’s L2RE to 1.017e-2. Comparing to the TB_Sigmoid, the L2RE decreased by 5.66%.
@@ 4.6 Ablation Study
Table 6: Different Window Size and Sampling Steps for pruning the LLaMA-7B at 0.8 sparsity ratio. The pruning method is SparseGPT. The model is evaluated on WikiText dataset with perplexity $(\downarrow)$.
@@ Appendix B Marchenko–Pastur Distribution
Figure 8: The Marchenko-Pastur (MP) Law for different values of $n$. The ESD (blue histogram) of the eigenvalues of the sample covariance matrix is compared with the theoretical Marchenko-Pastur probability density function (red curve) for various aspect ratios, where $m=4000$ is fixed, and $n$ varies from 4000 to 16000.
@@
Let $X_{ij},\,1\leq i\leq m,\,1\leq j\leq n$, be independent random variables with $\mathbb{E}X_{ij}=0$ and $\mathbb{E}X_{ij}^{2}=1$, and $\mathbf{X}_{m}=(X_{ij})_{1\leq i\leq m,1\leq j\leq n}$. Denote by $\lambda_{1}\leq\dots\leq\lambda_{m}$ the eigenvalues of the symmetric matrix [^4] Note. The transpose operation applies to the second (right-hand) matrix in this expression, which differs from the transpose placement used in HT-SR Theory, shown as $\mathbf{X}_{ij}=\mathbf{W}_{ij}^{\top}\mathbf{W}_{ij}$ in Section 3.3. Thus, the “large aspect ratio” regime discussed in HT-SR Theory corresponds to a large value of the ratio $n/m$ in Appendix B.
@@ C.2 Comparison of Different CNN Processing Methods
[…]
@@ C.3 Experiment Results for Additional Models
We prune the OPT models  (Zhang et al., 2022) and LLaMA-V2/V3 models  (Touvron et al., 2023b; Grattafiori et al., 2024) with AlphaPruning and compare the model performance with FARMS. We report the additional experiment results on LLM pruning in Table  8 and Table  9. For the OPT series models, we use four sparsity ratios: {0.6, 0.7, 0.75, 0.8}. For the LLaMA series models, we select {0.7, 0.75, 0.8, 0.85} as the sparsity ratios. Experimental results show that our method can help AlphaPruning achieve better model performance on models such as OPT.